rvc (rvc):

for x not = pi/2 is continous at x=pi/2 find f(pi/2)

rvc (rvc):

\[f(x)=\sqrt{2 }-\sqrt{1+\sin x}\div \cos ^2x\]

rvc (rvc):

@ikram002p

OpenStudy (ikram002p):

but its not continues at pi/2 , do u mean find limit?

rvc (rvc):

mmm yes

rvc (rvc):

@ikram002p

OpenStudy (ikram002p):

\(\Huge \frac {\sqrt 2 -\sqrt {1+\sin x} }{\cos^2 x}\) hint :- use LH rule do you know what it is ?

rvc (rvc):

no

rvc (rvc):

@ikram002p

OpenStudy (ikram002p):

ohh :) so if the limit was 0/0 or somthing/0 then use LH rule brb to explain

rvc (rvc):

@kropot72 pls help me

rvc (rvc):

@IMStuck

OpenStudy (ikram002p):

so to use LH rule to find the limits ,if u have a function of numerator and denominator u only have to find derivative of both numerator and denominator and then try to find the limit , its like this \(lim \dfrac {g(x)}{h(x)}=lim \dfrac {g'(x)}{h'(x)}\) \(lim \dfrac {g'(x)}{h'(x)}=lim \dfrac {g''(x)}{h''(x)}\)

rvc (rvc):

\[(\cos x \sin 2x)\div 2\sqrt{1+\sin x}\]

OpenStudy (ikram002p):

@dan815 could you continue here plz ? @rvc is my cute friend she is nice but i have to go for a while :'(

rvc (rvc):

@dan815

OpenStudy (dan815):

okya so lets see if this is still indeterminate form

OpenStudy (dan815):

lhopital can be reapplied as long as its indeterminate form, or now

rvc (rvc):

idk that

rvc (rvc):

give me the rule pls

OpenStudy (dan815):

okay so when its 0/0, at the given limit or +/- inf/inf at the given limit, we can differentiate the top and bottom of the fraction separately and it will equal the same thing

OpenStudy (dan815):

lets continue where u left off

OpenStudy (dan815):

\[\Huge \frac {\sqrt 2 -\sqrt {1+\sin x} }{\cos^2 x}\]

rvc (rvc):

okay

OpenStudy (dan815):

at f(pi/2) we have 0/0 so lets apply LH to this

rvc (rvc):

i have written the derivate

OpenStudy (dan815):

\[\Huge \frac {\sqrt 2 -\sqrt {1+\sin x} }{\cos^2 x}\] applying LH gives \[\frac{1/2\sqrt {(1+sinx)}*cosx}{2cosx*-sinx}\]

rvc (rvc):

have i done right?

OpenStudy (dan815):

nope because still indeterminate when evaluated at pi/2 so we can apply again

OpenStudy (dan815):

oh wait no need :)

OpenStudy (dan815):

u can cancel the cos and then it can be evaluated

rvc (rvc):

-2cosx*sinx=-sin2x

OpenStudy (dan815):

\[\frac{1/2\sqrt {(1+sinx)}}{2*-sinx}\] \[evaluate~ at~\pi/2 \]

OpenStudy (dan815):

\[\sqrt2/-4\]

rvc (rvc):

we can write sin2x

rvc (rvc):

?

rvc (rvc):

@dan815

OpenStudy (dan815):

that is not good to do because u will have a 0 in the denominator

OpenStudy (dan815):

hold on

rvc (rvc):

@dan815

OpenStudy (dan815):

okay yep it is right i believe

OpenStudy (dan815):

wait unless there is no negative

OpenStudy (dan815):

look at this

OpenStudy (dan815):

oh ya oops forgot a negative on top

OpenStudy (dan815):

it should be \[\sqrt2/4\]